OpenStudy (anonymous):

Recently a light powered by gravity has been gaining attention as an option for off-grid lighting. A falling weight pulls a cable to drive a generator, producing power for a LED light. (When the weight gets to the bottom, it must be lifted to begin the power generation again. In this sense it is a human powered light. Question: Assuming the human is 15% efficient in their use of food (chemical energy), how much chemical energy (in kJ) must be consumed by the human (i.e. ignoring basal metabolic rate) to power this light for 4 hours of operation? I'd appreciate any ideas. Thanks! :)

OpenStudy (anonymous):

@maqson :)

OpenStudy (aaronq):

How much energy does powering the light consume?

OpenStudy (anonymous):

the maximum average power that can be generated by letting the 5 kg weight fall the 1.5 meters back to the floor in 15 minutes is 0.0817 watts. is that what you mean?

OpenStudy (aaronq):

hm so it takes 15 min to fall to the ground?

OpenStudy (anonymous):

no, it doesnt say that. i gave all the information available. i think the idea is to assume that the weight doesnt reach the ground in the four hours.

OpenStudy (aaronq):

hm i think it does. So you have to calculate the \(\color{red}{energy~needed}\) to raise the 5 kg block 1.5 meters, every 15 min for 4 hrs, (i.e. raise the block 16 times). Then, assuming 15% efficiency: energy to be consumed(0.15)=\(\color{red}{energy~needed}\) energy to be consumed =\(\color{red}{energy~needed}\)/0.15

OpenStudy (anonymous):

The potential energy increase when lifted up 1.5 m is 73.5498752148 Joules. So altogether we have En = 16*73.5498752148 J = 1176.7980034368 J which gives Ec = 1176.7980034368/0.15 = 7845.3200 J = 7.8453 kJ. Thanks, great explanation!

OpenStudy (aaronq):

good stuff. no problem, dude !

OpenStudy (anonymous):

@techhelper hope you got the ans :)

OpenStudy (anonymous):

yes ;)